A ball is thrown straight up with a speed of 15m/s near the surface of the earth. What is the Maximum height of the ball?(neglect air friction)
1
Expert's answer
2020-10-23T11:53:12-0400

We can find the maximum height of the ball from the kinematic equation:


"v^2=v_0^2+2gh,"

here, "v=0\ \dfrac{m}{s}" is the speed of the ball at the maximum height, "v_0=15\ \dfrac{m}{s}" is the initial speed of the ball, "h" is the maximum height of the ball and "g=-9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Then, from this equation we can find the maximum height of the ball:


"h=\dfrac{-v_0^2}{2g}=\dfrac{-(15\ \dfrac{m}{s})^2}{2\cdot (-9.8\ \dfrac{m}{s^2})}=11.5\ m."

Answer:

"h=11.5\ m."